Statistics. Sam must have an average of 70 or more in his summer course to
obtain a grade of C. His first three test grades were 75, 63, and 68. how do I an inequality representing the score that Sam must get on the last test to get a C grade.

Answer by stanbon(48568)

Sam must have an average of 70 or more in his summer course to
obtain a grade of C. His first three test grades were 75, 63, and 68. how do I an inequality representing the score that Sam must get on the last test to get a C grade.
-----------------
Let the score on the last test be "x".
EQUATION:
[75+63+68+x}/4 = 70
You divide by 4 to get the average of the 4 test scores.
[75+63+68+x] = 280
206+x=280
x=74
He needs a score of 74 on the final test to get a C grade.
---------
Cheers,
Stan H.